I am having a hard time with this question. could somebody PLEASE!!!!!!
HELP ME.
Find the value of s if the perimeter of the square is 36 inches

this is the square
____________ (s+4)


____________

Answer by Paul(988)

SInce there are 4 equal sides in a square we represent the equation like this:
4(x+4)=36
4x+16=36
4x=36-16
4x=20
x=5
Hence the value of s is 5cm.